# The underlying structures of self-attention: symmetry, directionality, and emergent dynamics in Transformer training

### v1

@@ Acknowledgments
First, we show that a token $t^{*}$ contributes differently to the updates of $\mathbf{W}_{qk}$ depending on whether it serves as context for predicting other tokens or is itself predicted. When $t^{*}$ is used to predict a set of tokens:
  • All predicted tokens contribute to the column space of $\mathbf{W}_{qk}$ with the update of the $k$-th column scaled by the associated entry $[\mathbf{x}_{t^{*}}]_{k}$.
  • Only the token $t^{*}$ contributes to the row space, with the update to the $m$-th row scaled a linear combination of $[\mathbf{x}_{i}]_{m}$ for all predicted tokens.
On the other hand, when $t^{*}$ is predicted by a set of tokens:
  • Only the token $t^{*}$ contributes to the column space, with the update to the $k$-th row scaled a linear combination of $[\mathbf{x}_{i}]_{k}$ for all predicted tokens.
  • All context tokens contribute to the row space with the update of the $m$-th column scaled by the associated entry $[\mathbf{x}_{t^{*}}]_{m}$.
We formalize this in the following Proposition.
@@
(Different implicit updates for context and prediction). Let $U=[t_{0},\dots,t_{N}]$ be a sequence of tokens and let $\mathbf{x}_{i}\in\mathbb{R}^{d}$ be the embedding of the $i$-th token. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.2. Let $t^{*}$ be a given token in $U$. When using $t^{*}$ as context, the $k$-th column of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{\cdot,k}\bigg{|}_{t_{j}=t^{*}}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in P_{t^{*}}}\beta_{ij}\mathbf{x}_{i}\right)\,,$ (S42)
while the $m$-th row of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in P_{t^{*}}}\beta_{ij}[\mathbf{x}_{i}]_{m}\right)\mathbf{x}_{t^{*}}\,,$ (S43)
where $P_{t^{*}}$ is the set of tokens predicted by $t^{*}$. When predicting $t^{*}$, the $k$-th column of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{k,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}[\mathbf{x}_{j}]_{k}\right)\mathbf{x}_{t^{*}}\,,$ (S44)
while the $m$-th row of $\Delta\mathbf{W}_{qk}$ is given by
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{i}=t^{*}}=[\mathbf{x}_{t^{*}}]_{m}\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}\mathbf{x}_{j}^{\top}\right)\,.$ (S45)
where $C_{t^{*}}$ is the set of context tokens for $t^{*}$.
@@
Let $\mathbf{U}=[t_{0},\dots,t_{N}]$ a sequence of tokens with the embedding of every $i$-th token be given by $\mathbf{x}_{i}\in\mathbb{R}^{d}$. The implicit weight update of Proposition 2.2 can be decomposed with two equivalent regrouping of the double summations, as follows,
$\Delta\mathbf{W}_{qk}=\sum_{(i,j)\in\mathbf{U}}\beta_{ij}\mathbf{x}_{i}\mathbf{x}_{j}^{T}=\sum_{j}\left(\sum_{i\in P_{j}}\beta_{ij}\mathbf{x}_{i}\right)\mathbf{x}^{\top}_{j}=\sum_{i}\mathbf{x}_{i}\left(\sum_{i\in C_{i}}\beta_{ij}\mathbf{x}^{\top}_{j}\right)\,,$ (S46)
where $P_{i}\subset U$ is the set of tokens predicted by a given token $t_{i}$, while $C_{j}\subset U$ is the set of tokens that predict a given token $t_{j}$. We neglect any constant of proportionality - such as a learning rate - for simplicity. For simplicity, we do not assume any specific structure on $P_{i}$ and $C_{j}$ (autoregressive training, bidirectional training, or others). First, the contribution of $t^{*}\in U$ to the weight update when $t^{*}$ is used as context to predict a set of tokens $P_{t^{*}}\subset U$ is
$\Delta\mathbf{W}_{qk}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in P_{t^{*}}}\beta_{it^{*}}\mathbf{x}_{i}\right)\mathbf{x}^{\top}_{t^{*}}\,.$ (S47)
The associated weight update of the $k$-th column is then given by
@@
$\Delta\mathbf{w}_{\cdot,k}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in P_{t^{*}}}\beta_{it^{*}}\mathbf{x}_{i}\right)[\mathbf{x}_{t^{*}}]_{k}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in P_{t^{*}}}\beta_{it^{*}}\mathbf{x}_{i}\right)\,,$ (S48)
while the update of the $m$-th row is
@@
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in P_{t^{*}}}\beta_{it^{*}}[\mathbf{x}_{i}]_{m}\right)\mathbf{x}_{t^{*}}\,.$ (S49)
A complementary argument can be made when predicting a given token $t^{*}$ from a set of tokens $C_{t^{*}}$. Indeed, the contribution to the weight update when $t^{*}$ is predicted by set of tokens $C_{t^{*}}\subset U$ is given by
$\Delta\mathbf{W}_{qk}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}\mathbf{x}_{t^{*}}\right)\mathbf{x}_{j}^{\top}\,.$ (S50)
The associated weight update of the $k$-th column is then given by
@@
$\Delta\mathbf{w}_{k,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}[\mathbf{x}_{j}]_{k}\right)\mathbf{x}_{t^{*}}\,,$ (S51)
while the update of the $m$-th row is
@@
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}\mathbf{x}_{j}^{\top}\right)[\mathbf{x}_{t^{*}}]_{m}=[\mathbf{x}_{t^{*}}]_{m}\left(\sum_{j\in C_{t^{*}}}\beta_{t^{*}j}\mathbf{x}_{j}^{\top}\right)\,.$ (S52)
S1.4.2 Asymmetric growth of rows and columns during weight update
Next, we demonstrate that under reasonable assumptions about the statistical distribution of token embeddings, the expected norm of column updates exceeds that of row updates when using the token $t^{*}$ to predict other tokens. Conversely, when $t^{*}$ is being predicted by other tokens, the row updates become dominant. We begin by assuming that the token embeddings $\mathbf{x}_{i}$ are independent and identically distributed (i.i.d.) random vectors drawn from a probability distribution $\mathcal{D}$ with zero mean, $\mathbb{E}[\mathbf{x}_{i}]=\mathbf{0}$, and covariance matrix $\text{Cov}(\mathbf{x}_{i})=\Sigma$. This assumption holds true at initialization for any Transformer model with learnable embeddings. Additionally, we assume that the covariance matrix satisfies $\Sigma\neq\sigma^{2}\mathbb{I}$,. More specifically, we posit that there is partial alignment between the embeddings $\mathbf{x}_{i}$ due to the semantic and predictive relationships between tokens, which typically emerge in the vector embeddings during training. Similar to Proposition S1.7, the scenarios where $t^{*}$ is used to predict other tokens and where $t^{*}$ is being predicted by other tokens are complementary. In the following Proposition, we focus solely on the case where $t^{*}$ serves as context to predict a set of tokens. A formal derivation for the opposite case where $t^{*}$ is predicted by other tokens is provided in a subsequent Corollary.
@@
(Asymmetric growth of columns and rows for context). Let $U=[t_{0},\dots,t_{N}]$ a sequence of tokens, and let $t^{*}$ be a token representing the context of every token $t_{i}\in U$. Let $\{\mathbf{x}_{1}\dots,\mathbf{x}_{N}\}$ be the token embedding associated with $U$ such that each $\mathbf{x}_{i}\sim\mathcal{D}$ is a i.i.d. random vector drawn from a probability distribution $\mathcal{D}$ with zero mean and non-isotropic covariance $\text{Cov}(\mathbf{x}_{i})=\Sigma$. Let $\mathbf{y}$ be the independent random embedding of $t^{*}$ with zero mean and covariance $\text{Cov}(\mathbf{y})=\sigma_{y}^{2}\mathbb{I}$. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.2. Then the squared norm of the $m$-th row and $k$-th column of $\mathbf{W}_{qk}$ satisfies
$\frac{\mathbb{E}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}>1\quad\forall k\in\{1,\dots,d\},\forall m\,\,\text{s.t.}\,\,\Sigma_{m,m}<\frac{\text{Tr}(\Sigma)}{d}$ (S53)

### v2

@@ Acknowledgments
First, we show that a token $t^{*}$ contributes differently to the updates of $\mathbf{W}_{qk}$ depending on whether it serves as context for predicting other tokens or is itself predicted.
Proof.
Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of tokens with the embedding of every $i$-th token be given by $\mathbf{x}_{i}\in\mathbb{R}^{d}$. Proposition 2.1 shows that the implicit weight update can be decomposed with two equivalent regrouping of the double summations, as follows,
$\Delta\mathbf{W}_{qk}=\sum_{(i,j)\in U}\beta_{ij}\mathbf{x}_{i}\mathbf{x}_{j}^{T}=\sum_{j}\left(\sum_{i\in\mathcal{P}_{j}}\beta_{ij}\mathbf{x}_{i}\right)\mathbf{x}^{\top}_{j}=\sum_{i}\mathbf{x}_{i}\left(\sum_{i\in\mathcal{C}_{i}}\beta_{ij}\mathbf{x}^{\top}_{j}\right)\,,$ (S40)
where $\mathcal{P}_{i}\subset U$ is the set of tokens predicted by a given token $t_{i}$, while $\mathcal{C}_{j}\subset U$ is the set of tokens that predict a given token $t_{j}$. We neglect any constant of proportionality - such as a learning rate - for simplicity, and we do not assume any specific structure on $\mathcal{P}_{i}$ and $\mathcal{C}_{j}$ (autoregressive training, bidirectional training, or others). First, the contribution of $t^{*}\in U$ to the weight update when $t^{*}$ is used as context to predict a set of tokens $\mathcal{P}_{t^{*}}\subset U$ is
$\Delta\mathbf{W}_{qk}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}\mathbf{x}_{i}\right)\mathbf{x}^{\top}_{t^{*}}\,.$ (S41)
The associated weight update of the $k$-th column is then given by
$\Delta\mathbf{w}_{\cdot,k}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}\mathbf{x}_{i}\right)[\mathbf{x}_{t^{*}}]_{k}=[\mathbf{x}_{t^{*}}]_{k}\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}\mathbf{x}_{i}\right)\,,$ (S42)
while the update of the $m$-th row is
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{j}=t^{*}}=\left(\sum_{i\in\mathcal{P}_{t^{*}}}\beta_{i*}[\mathbf{x}_{i}]_{m}\right)\mathbf{x}_{t^{*}}\,.$ (S43)
A complementary argument can be made when predicting a given token $t^{*}$ from a set of tokens $\mathcal{C}_{t^{*}}$. Indeed, the contribution to the weight update when $t^{*}$ is predicted by set of tokens $\mathcal{C}_{t^{*}}\subset U$ is given by
$\Delta\mathbf{W}_{qk}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}\mathbf{x}_{t^{*}}\right)\mathbf{x}_{j}^{\top}\,.$ (S44)
The associated weight update of the $k$-th column is then given by
$\Delta\mathbf{w}_{k,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}[\mathbf{x}_{j}]_{k}\right)\mathbf{x}_{t^{*}}\,,$ (S45)
while the update of the $m$-th row is
$\Delta\mathbf{w}_{m,\cdot}\bigg{|}_{t_{i}=t^{*}}=\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}\mathbf{x}_{j}^{\top}\right)[\mathbf{x}_{t^{*}}]_{m}=[\mathbf{x}_{t^{*}}]_{m}\left(\sum_{j\in\mathcal{C}_{t^{*}}}\beta_{*j}\mathbf{x}_{j}^{\top}\right)\,.$ (S46)
S1.5.2 Asymmetric growth of rows and columns during weight update
Next, we demonstrate that under reasonable assumptions about the statistical distribution of token embeddings, the expected norm of column updates exceeds that of row updates when the token $t^{*}$ is used as context to predict other tokens. Conversely, when $t^{*}$ is being predicted by other tokens, the row updates become dominant. To do so, we make the following assumptions:
  • The token embeddings $\mathbf{x}_{i}$ are independent and identically distributed (i.i.d.) random vectors drawn from a probability distribution $\mathcal{D}$ with zero mean, $\mathbb{E}[\mathbf{x}_{i}]=\mathbf{0}$, and covariance matrix $\text{Cov}(\mathbf{x}_{i})=\Sigma$. This assumption holds at initialization for any Transformer model with learnable embeddings.
  • The covariance matrix is not isotropic, i.e., $\Sigma\neq\sigma^{2}\mathbb{I}$,. More specifically, we posit that there is partial alignment between the embeddings $\mathbf{x}_{i}$ due to the semantic and predictive relationships between tokens, which typically emerge during training.
Similar to Proposition 2.2, the scenarios where $t^{*}$ is used to predict other tokens and where $t^{*}$ is being predicted by other tokens are complementary. In the following Proposition, we focus solely on the case where $t^{*}$ serves as context to predict a set of tokens. A formal derivation for the opposite case where $t^{*}$ is predicted by other tokens is provided in a subsequent Corollary.
@@
(Asymmetric growth of columns and rows for context). Let $U=\{t_{1},\dots,t_{N}\}$ a sequence of tokens, and let $t^{*}$ be a token representing the context of every token $t_{i}\in U$. Let $\{\mathbf{x}_{1}\dots,\mathbf{x}_{N}\}$ be the token embedding associated with $U$ such that each $\mathbf{x}_{i}\sim\mathcal{D}$ is a i.i.d. random vector drawn from a probability distribution $\mathcal{D}$ with zero mean and non-isotropic covariance $\text{Cov}(\mathbf{x}_{i})=\Sigma$. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.1. Then the squared norm of the $m$-th row and $k$-th column of $\mathbf{W}_{qk}$ satisfies
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}>1\quad\forall k\in\{1,\dots,d\},\forall m\,\,\text{s.t.}\,\,\Sigma_{m,m}<\frac{\text{Tr}(\Sigma)}{d}$ (S47)
@@
Let $U=[t_{0},\dots,t_{N}]$ be a sequence of tokens, and let $t^{*}$ be the context of every token $t_{i}\in U$. Let $\{\mathbf{x}_{i}\}$ be the set of token embeddings associated with $U$ and let $\mathbf{y}$ be the token embedding of $t^{*}$. It follows from Proposition 2.2 that the squared norm of the weight update of the $k$-th column is given by
$\Delta\mathbf{w}_{\cdot,k}=\sum_{i=1}^{N}\beta_{i*}\mathbf{x}_{i}[\mathbf{y}]_{k}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{\cdot,k}||^{2}=[\mathbf{y}]^{2}_{k}||\sum_{i=1}^{N}\beta_{i*}\mathbf{x}_{i}||^{2}\,,$ (S48)
while the squared norm of the weight update of the $m$-th row by
$\Delta\mathbf{w}_{m,\cdot}=\sum_{i=1}^{N}\beta_{i*}[\mathbf{x}_{i}]_{m}\mathbf{y}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{m,\cdot}||^{2}=\big{(}\sum_{i=1}^{N}\beta_{i*}[\mathbf{x}_{i}]_{m}\big{)}^{2}||\mathbf{y}||^{2}\,.$ (S49)
Let $\{\mathbf{x}_{i}\}\sim\mathcal{D}$ be a set of i.i.d. random vectors from a probability distribution $\mathcal{D}$ such that $\mathbb{E}_{\mathcal{D}}[\mathbf{x}_{i}]=0$ and $\text{Cov}(\mathbf{x}_{i})=\Sigma\in\mathbb{R}^{d,d}$. Therefore, each $k$-th coordinate $[\mathbf{x}_{i}]_{k}$ is such that $\mathbb{E}_{\mathcal{D}}[[\mathbf{x}_{i}]_{k}]=0$ and $\text{Var}([\mathbf{x}_{i}]_{k})=\Sigma_{k,k}$. We also assume that $\mathbf{y}$ is statistically independent from $\mathbf{x}_{i}\,\forall i$ with $\mathbb{E}[\mathbf{y}]=0$ and covariance $\text{Cov}(\mathbf{y})=\sigma_{y}^{2}\mathbb{I}$. It follows that the expected value of $||\Delta\mathbf{w}_{\cdot,k}||^{2}$ over $\mathcal{D}$ is given by
@@
$\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]=\mathbb{E}_{\mathcal{D}}\left[[\mathbf{y}]^{2}_{k}||\sum_{i=1}^{N}\beta_{i*}\mathbf{x}_{i}||^{2}\right]=\mathbb{E}_{\mathcal{D}}\left[[\mathbf{y}]^{2}_{k}\right]\mathbb{E}_{\mathcal{D}}\left[||\sum_{i=1}^{N}\beta_{i*}\mathbf{x}_{i}||^{2}\right]\,.$ (S50)
$\begin{split}\mathbb{E}_{\mathcal{D}}\left[||\sum_{i=1}^{N}\beta_{i*}\mathbf{x}_{i}||^{2}\right]&=\sum_{i,i^{\prime}}\mathbb{E}_{\mathcal{D}}\left[\beta_{i*}\beta_{i^{\prime}*}\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\sum_{i=1}^{N}\beta^{2}_{i*}\mathbb{E}_{\mathcal{D}}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i}\right]+\sum_{i^{\prime}\neq i}\beta_{i*}\beta_{i^{\prime}*}\mathbb{E}_{\mathcal{D}}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\sum_{i=1}^{N}\beta^{2}_{i*}\text{Tr}\big{(}\mathbb{E}_{\mathcal{D}}\left[\mathbf{x}_{i}\mathbf{x}_{i}^{\top}\right]\big{)}+\sum_{i^{\prime}\neq i}\beta_{i*}\beta_{i^{\prime}*}\mathbb{E}_{\mathcal{D}}\left[\mathbf{x}^{\top}_{i}\mathbf{x}_{i^{\prime}}\right]\\ &=\text{Tr}(\Sigma)\sum_{i=1}^{N}\beta^{2}_{i*}\,,\end{split}$ (S51)
and therefore
@@
$\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]=\Gamma_{k,k}\text{Tr}(\Sigma)\sum_{i=1}^{N}\beta^{2}_{i*}\,.$ (S52)
@@
$\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]=\mathbb{E}_{\mathcal{D}}\left[\big{(}\sum_{i=1}^{N}\beta_{i*}[\mathbf{x}_{i}]_{m}\big{)}^{2}||\mathbf{y}||^{2}\right]=\mathbb{E}_{\mathcal{D}}\left[\big{(}\sum_{i=1}^{N}\beta_{i*}[\mathbf{x}_{i}]_{m}\big{)}^{2}\right]\mathbb{E}_{\mathcal{D}}\left[||\mathbf{y}||^{2}\right]\,.$ (S53)
The first term can be decomposed as
$\mathbb{E}_{\mathcal{D}}\left[\big{(}\sum_{i=1}^{N}\beta_{i*}[\mathbf{x}_{i}]_{m}\big{)}^{2}\right]=\sum_{i=1}^{N}\beta^{2}_{i*}\mathbb{E}_{\mathcal{D}}\left[[\mathbf{x}_{i}]_{m}^{2}\right]+\sum_{i^{\prime}\neq i}\beta_{i*}\beta_{i^{\prime}*}\mathbb{E}_{\mathcal{D}}\left[[\mathbf{x}_{i^{\prime}}]_{m}[\mathbf{x}_{i}]_{m}\right]=\Sigma_{m,m}\sum_{i=1}^{N}\beta^{2}_{i*}\,,$ (S54)
and therefore
@@
$\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]=\Sigma_{m,m}\text{Tr}(\Gamma)\sum_{i=1}^{N}\beta^{2}_{i*}\,.$ (S55)
The ratio of the expected value of these squared norms is
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{\Gamma_{k,k}\text{Tr}(\Sigma)}{\Sigma_{m,m}\text{Tr}(\Gamma)}=\frac{\Gamma_{k,k}}{\text{Tr}(\Gamma)}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}=\frac{1}{d}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}\,.$ (S56)
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{1}{d}\frac{\text{Tr}(\Sigma)}{\Sigma_{m,m}}>\frac{1}{d}\frac{d\,\Sigma_{m,m}}{\Sigma_{m,m}}=1\quad\forall m\,\,\text{s.t.}\,\,\Sigma_{m,m}<\frac{\text{Tr}(\Sigma)}{d}\,,$ (S57)
thus concluding the proof. ∎
Again, a complementary argument can be made about the asymmetric weight update when a set of tokens $U=\{t_{1},\dots,t_{N}\}$ is used to predict a given token $t^{*}$. We formalize this in the following Corollary.
Corollary S1.8.
(Asymmetric growth of columns and rows for prediction). Let $U=\{t_{1},\dots,t_{N}\}$ be a sequence of tokens, and let every token $t_{i}\in U$ be the context of a given token $t^{*}$. Let the associated embedding be drawn i.i.d. from a probability distribution $\mathcal{D}$ as in Proposition S1.7. Let $\Delta\mathbf{W}_{qk}$ be the weight update from Proposition 2.1. Then the squared norm of the $m$-th row and $k$-th column of $\mathbf{W}_{qk}$ satisfies
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}<1\quad\forall m\in\{1,\dots,d\},\forall k\,\,\text{s.t.}\,\,\Sigma_{k,k}<\frac{\text{Tr}(\Sigma)}{d}$ (S58)
Proof.
Let $U=[t_{0},\dots,t_{N}]$ a sequence of tokens, and let $t^{*}$ be predicted by every token $t_{i}\in U$. Let $\{\mathbf{x}_{i}\}$ be the set of token embeddings associated with $U$ and let $\mathbf{y}$ be the token embedding of $t^{*}$. It follows from Proposition 2.2 that the squared norm of the weight update of the $k$-th column is given by
$\Delta\mathbf{w}_{\cdot,k}=\sum_{j=1}^{N}\beta_{*j}[\mathbf{x}_{j}]_{k}\mathbf{y}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{\cdot,k}||^{2}=\big{(}\sum_{i=1}^{N}\beta_{*j}[\mathbf{x}_{j}]_{k}\big{)}^{2}||\mathbf{y}||^{2}\,,$ (S59)
while the weight update of the $m$-th row follows
$\Delta\mathbf{w}_{m,\cdot}=\sum_{j=1}^{N}\beta_{*j}\mathbf{x}_{j}[\mathbf{y}]_{m}\,\,\Rightarrow\,\,||\Delta\mathbf{w}_{m,\cdot}||^{2}=[\mathbf{y}]^{2}_{m}||\sum_{i=1}^{N}\beta_{*j}\mathbf{x}_{j}||^{2}\,.$ (S60)
Therefore, following the same arguments as in Proposition S1.7, the ratio of the expected value of the square norms is given by
$\frac{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{\cdot,k}||^{2}\right]}{\mathbb{E}_{\mathcal{D}}\left[||\Delta\mathbf{w}_{m,\cdot}||^{2}\right]}=\frac{\Sigma_{k,k}\text{Tr}(\Gamma)}{\Gamma_{m,m}\text{Tr}(\Sigma)}=d\frac{\Sigma_{k,k}}{\text{Tr}(\Sigma)}<\frac{d\,\Sigma_{k,k}}{d\,\Sigma_{k,k}}=1\quad\forall k\,\,\text{s.t.}\,\,\Sigma_{k,k}<\frac{\text{Tr}(\Sigma)}{d}\,,$ (S61)
thus concluding the proof. ∎
S1.5.3 Expected contribution for context and prediction and related remarks
Additionally, we show that in autoregressive training, the expected number of tokens predicted by a given token $t^{*}$ can differ from the number of tokens that predict $t^{*}$, with this ratio influenced by token correlations, while in bidirectional training, the two quantities are always equal regardless of such correlations, yielding a ratio of 1. We formalize this in the following proposition and illustrate, in the subsequent remark, an example from autoregressive training where the expected ratio exceeds 1.
@@
(Ratio of expected counts during autoregressive and bidirectional training). Let $\mathcal{V}=\{t_{0},\dots,t_{V}\}$ be a set of tokens. Let $\mathcal{U}$ be the sample space of all possible sequences of $N$ tokens, and let the sequence $U=\{t_{1},\dots,t_{N}\}\in\mathcal{U}$ be a random variable with probability distribution $P(U)$ defined over $\mathcal{U}$. Let $\Pr[t_{j}=t^{*}]$ be the probability that the token at index $j$ in a sequence $U\in\mathcal{U}$ is given by $t^{*}\in\mathcal{V}$. Let $\mathbb{E}_{P}[\mu_{c}(t^{*})]$ be the expected number of tokens that are predicted by a given token $t^{*}$, and $\mathbb{E}_{P}[\mu_{p}(t^{*})]$ the expected number of tokens that predict a given token $t^{*}$. For autoregressive training, the ratio
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=\frac{\sum_{k=1}^{N}(N-k)\Pr[t_{k}=t^{*}]}{\sum_{k=1}^{N}(k-1)\Pr[t_{k}=t^{*}]}\,,$ (S62)
depends on $\Pr[t_{k}=t^{*}]\,\,\forall t^{*}\in\mathcal{V},\forall k\in\{1,\dots.N\}$, while for bidirectional training the same ratio is given by
$\frac{\mathbb{E}_{P}[\mu_{c}(t^{*},U)]}{\mathbb{E}_{P}[\mu_{p}(t^{*},U)]}=1\,.$ (S63)